a lawn ahs dimensions 8m by 8m. how wide a strip must be mowed around the outside edge so that exactly half of the lawn is cut? Give your answer in simplest radical for, use the square root property to solve the equation that results.
Answer by venugopalramana(3286)

a lawn ahs dimensions 8m by 8m. how wide a strip must be mowed around the outside edge so that exactly half of the lawn is cut? Give your answer in simplest radical for, use the square root property to solve the equation that results.
AREA OF LAWN =8*8=64
HALF OF IT =64/2=32=AREA OF STRIP TO BE REMOVED.IF THE STRIP IS OF SIZE X*X,THEN AREA OF STRIPCUT IS = 64-X^2=32
X^2=32
X=SQRT(32)=4SQRT(2)
SO IF WIDTH OF STRIPS ON EITHER SIDE IS W ,THEN TOTAL WIDTH OF STRIPS =2W =(8-X)
SO....W=(8-X)/2=(8-4SQRT(2))/2 =4-2SQRT(2)